OpenStudy (anonymous):

Find the slope of the line whose equation is 5y = x - 3. -3 -3/5 1/5 5

OpenStudy (anonymous):

There are several ways to answer this question. The choice that is easiest depends on the form the equation is in ! If it is in slope intercept form, which is y=mx +b, the slope is the coefficient of the x term! Be sure the coefficient of the y term is one, or it is not in slope intercept form ! If the equation is in standard form, which is Ax + By = C , basically the x and y terms are on one side of the equation and the constant is on the other, then the slope is -A/B . You could also find two points on the line and use the slope formula !!

OpenStudy (anonymous):

So, 5y = x-3 becomes y = 1/5 x - 3/5 , so the slope is 1/5. Or, put the equation in the form -x + 5y = -3 so, slope is -A/B or -(-1) /5 which equals 1/5.